FOURIER-TRANSFORM SPECTROSCOPY OF ^{14}NH_{3} AND ^{15}NH_{3} IN THE NEAR-INFRARED

N. IBRAHIM, P. CHELIN, and J. ORPHAL, Laboratoire Interuniversitaire des Systèmes Atmosphériques (LISA), CNRS and Universités Paris-7 et -12, Créteil, France; C. E. FELLOWS, Laboratório de Espectroscopia e Laser, Instituto de Física, Universidade Federal Fluminense, Niterói, Rio de Janeiro, Brazil.

The near-infrared absorption spectrum of NH_{3} is an interesting opportunity for atmospheric in-situ measurements of this species using tuneable diode-lasers. For this purpose, accurate line positions and intensities are indispensable. In the past, several studies of the absorption spectrum of ^{14}NH_{3} in the near-infrared were carried out, using Fouriertransform^{1} and tuneable diode-laser spectroscopy.^{2−4} More recently, high-resolution diode-laser spectra of ^{15}NH_{3} were analyzed for the first time.^{5} The lines of ^{14}NH_{3} in this region are also useful for wavenumber calibration of absorption spectra obtained with tuneable lasers. In our group, we have recently investigated the possibility to detect atmospheric NH_{3} using photoacoustic laser spectroscopy in the 1.5 μm region, and observed several discrepancies between individual line positions and intensities observed in our and previous studies^{1,4}, as already noticed by other groups.^{2,5} In order to solve this problem we have measured new absorption spectra of ^{14}NH_{3} (and also of ^{15}NH_{3}) in the 6300-7500 cm^{−1} region using a Bruker IFS 120-HR Fouriertransform spectrometer. The length of the absorption cell was 30 cm. The NH_{3} pressures employed were around 30 mbar, and the spectral resolution used was 0.02 cm^{−1}, leading to linewidths (Full-Width at Half-Maximum) of about 0.03 cm^{−1}. Based on these spectra, we have produced a list of individual line positions and line intensities at 296 K. Comparisons with the previous studies^{1−5} will be presented.

[1] L. Lundsberg-Nielsen et al., J. Mol. Spectrosc., 162, 230-245 (1993).

[2] M. Webber et al., Appl. Opt., 40, 2031-2042 (2004).

[3] L.-H. Xu et al., Infrared Phys. Technol., 45, 31-45 (2004).

[4] J. S. Gibb et al., Eur. Phys. J. D., 28, 59-66 (2004).

[5] R. M. Lees et al., J. Mol. Struct., in press (2006).